OpenStudy (anonymous):

Which point lies on the line 2x - y = 8?

OpenStudy (saifoo.khan):

as u r going to solve the equation

OpenStudy (saifoo.khan):

2x - y = 8 2x - 8 = y

OpenStudy (saifoo.khan):

Now, y = mx + b compare y = 2x - 8 slope is 2, whereas the y intercept is -8.